# A Comprehensive Survey of Small Language Models in the Era of Large Language Models: Techniques, Enhancements, Applications, Collaboration with LLMs, and Trustworthiness

### v1

@@ 4.1.5. SLM Applications in Mobile-device
The use of cloud-based LLMs on devices has raised privacy concerns, with their large size limiting practicality on mobile devices and hindering real-time responses in urgent scenarios such as medical emergencies. To address these limitations, researchers are developing smaller, domain-specific models that deliver accurate results, making SLMs a viable alternative. This subsection introduces SLM applications on mobile devices, categorizing works into three aspects: software API calls, mobile control, and basic NLP applications on devices.
SLM for Software API Call. Integrating LLMs with external APIs enhances their capabilities by accessing up-to-date information and specialized functionalities, but fine-tuning them for API calls incurs significant training costs. Consequently, there is a trend toward developing smaller, task-specific language models that preserve essential functionality while reducing training costs. Nonetheless, these smaller models pose increased risks of errors and precision issues, which are critical for solving various tasks. In response, Octopus (Chen et al., 2024a) crafts a diverse dataset from over 30,000 widely-used APIs and implements curriculum learning strategies (Liu et al., 2024e) to enhance accuracy in selecting the appropriate API functions. This approach has been applied in models such as Codellama-7b (Roziere et al., 2023) and Google’s Gemma series (Team et al., 2024a), significantly improving API call performance. For instance, Octopus-gemma2B achieves a function accuracy of 93% compared to ChatGPT’s 50% and closely approaches GPT-4’s 96% on their crafted dataset.
@@
SLM for Mobile Control. LLM agents facilitate user-device interactions through taps, gestures, and text, automating tasks and enhancing user hands-free convenience. Unlike traditional developer-based approaches that require extensive developer effort to design interfaces and translate commands into API calls, LLMs offer scalable automation via GUI-based text contents. MobileAgent (Ding, 2024) integrates instructions and Standard Operating Procedures (SOP) to enhance SLM performance in mobile control applications. An example, as shown in Figure 14, begins with the AI agent receiving a goal, such as booking a dental appointment. It analyzes the mobile screen and query, processes prior actions and environmental factors like UI elements, and forms a prompt. The agent then generates an output—choosing an appointment—and executes actions by identifying relevant UI elements on the mobile interface, such as text and XPath. They fine-tune the Qwen-7B model (Bai et al., 2023) on a dataset from the AIA medical application, evaluating it on the AitW benchmark (Rawles et al., 2024), where it demonstrates superior performance to GPT-4 (Achiam et al., 2023) without additional inference costs. AitW is recognized for its relevance in mobile operations. Carreira et al. (2023) addresses privacy and latency concerns of LLMs by running a small model offline on mobile devices. This model, fine-tuned with data generated by ChatGPT-3.5, achieves text-to-action capabilities for tasks such as making calls and conducting web searches. The RedPajama-INCITE-Chat-3B-v1  [^1] https://huggingface.co/togethercomputer/RedPajama-INCITE-Chat-3B-v1 model is chosen for its size and chatting capabilities, employing native code and model quantization techniques. Despite hardware constraints, it delivers adequate performance on 6GB Android devices and operates on 4GB devices.
@@
AutoDroid (Wen et al., 2024) is developed to enhance Android app interactions through a GUI. Figure 15 shows an example of LLM-powered mobile task automation (Carreira et al., 2023) that a user asking to be reminded about doing laundry on Aug 17. The process involves four steps: (1) clicking ’New Event’, (2) entering ’laundry’ in the ’Title’ field, (3) clicking ’Save’, and (4) completing the task. The agent uses the phone GUI, guided by Vicuna-7B and the app’s domain knowledge, to complete the task. AutoDroid dynamically integrates SLM-derived commonsense and app-specific knowledge, using it to generate privacy-filtered prompts that drive app interactions based on user commands. Its effectiveness is validated on the proposed DroidTask benchmark, showcasing superior performance over GPT-3.5 and GPT-4 (Achiam et al., 2023). For example, their Vicuna-7B AutoDroid achieves 57.7%, outperforming ChatGPT’s 34.7% and GPT-4’s 54.5% (Achiam et al., 2023). These studies illustrate the potential of customizing smaller, domain-specific SLMs in mobile environments, addressing traditional memory limitations while maintaining essential functionalities.
@@
Figure 16. Overview of the framework for fine-tuning LLMs using synthesized and user-generated data (Qin et al., 2023). Fine-tune LLMs using data from data selection and following data generation. The process involves Data Synthesis, where prompts are used to generate semantically similar text, creating new dialogue sets. Data Selection follows, where user-generated data is processed, and tagged with domain labels, and metrics such as EOE, DSS, and IDD are calculated. The selected data is used to fine-tune the LLM, with user annotations guiding the process. The framework continuously refines LLMs based on this iterative data generation and selection.
@@
SLM for Basic NLP Applications on Devices To enable personalization on mobile devices while ensuring privacy, performing basic NLP tasks such as text rewriting directly on the device is essential, avoiding data transfer to the cloud. Qin et al. (2023) introduces a framework that utilizes self-supervised data selection and synthesis for on-device fine-tuning, leveraging sparse annotations and limited storage effectively. This approach, demonstrated in Figure 16, employs the Llama-3B model (Touvron et al., 2023a) and the LoRA fine-tuning method (Hu et al., 2021), enhancing personalization by efficiently managing data through metrics like embedding entropy and domain-specific scores. In addressing mobile text rewriting, Zhu et al. (2023b) train the compact Palm 2-XXS model (Anil et al., 2023) using data generated by the larger Palm 2-L to ensure user privacy and accommodate device constraints. Their new evaluation benchmark, MESSAGEREWRITEEVAL, demonstrates that this mobile model surpasses the performance of traditional LLMs such as LLaMA (Touvron et al., 2023a) and Alpaca-7B (Taori et al., 2023). For example, the BLEU score of the proposed SFT Palm 2-XXS is 34.59, outperforming the 16.65 achieved by LLaMa 7B (Touvron et al., 2023a). Performance benchmarks on devices like the Samsung S23 and Pixel 7 Pro show significantly lower latency (36.2ms and 59.8ms, respectively) compared to tests conducted on a MacBook M1 Pro with a 4-bit quantized Llama 7B model (18-22 tokens/second), validating the effectiveness and practicality of this mobile-centric approach in text rewriting tasks.
@@
Insights: We draw several key insights from the development of task-specific SLMs: There is considerable potential in enhancing the efficiency and effectiveness of small models by integrating self-adaptive techniques with further fine-tuning and optimization of RAG-based methods. The growing relevance of SLMs in coding highlights their cost-effectiveness and efficiency as alternatives to LLMs, providing quick processing and easy fine-tuning for specialized tasks; while LLMs handle complex tasks well, SLMs, optimized and fine-tuned on specific data, are increasingly essential in resource-limited settings. SLMs significantly enhance recommendation systems due to their robust generalization, reasoning abilities, and in-context learning, addressing key challenges such as cold-start problems and distribution biases. They support long-term planning, replace traditional encoders, and use parallel low-rank parameters to inject personalized user knowledge effectively. SLMs play a crucial role in web search such as document encoding, text reordering, and query rewriting, often outperforming LLMs through techniques such as supervised fine-tuning, soft prompt tuning, unsupervised contrastive loss, and reinforcement learning, thereby enhancing adaptability and efficiency.
SLMs are utilized on mobile devices primarily for privacy and memory constraints, with applications in API calls and mobile control; they are typically developed by generating data with LLMs and fine-tuning with SLMs, or by using local SLMs to handle privacy with LLMs boosting performance, and their training involves innovative techniques like learning from data streams and managing non-IID time series data.
@@
AWQ (Lin et al., 2024c) Preserve key weights based on activation distribution, not weight distribution
@@
Transformer-Lite (Li et al., 2024f) Optimize KV cache to reduce redundancy and memory use.
@@
Runtime Efficiency Optimization EdgeMoE (Yi et al., 2023) Predict expert needs, boosting inference speed and reducing latency.
@@
4.2. SLM Deployment on Mobile and Edge Devices
4.2.1. Memory Efficiency Optimization
Quantization. Quantization, a prevalent and efficient approach in designing and deploying SLMs, reduces the model’s numerical precision to significantly save memory while maintaining accuracy. We outline quantization strategies in Sections 2.3.3 and 3.5. Here we focus on representative quantization works related to edge devices and their evaluations of memory usage. Regarding quantization works related to edge devices, the EDGE-LLM framework (Yu et al., 2024c) adapts LLMs for edge devices using a Layer-wise Unified Compression (LUC) method that combines layer-specific pruning and quantization to reduce computational demands and an Adaptive Layer Tuning and Voting scheme to optimize memory use while ensuring performance. Meanwhile, LLM-PQ (Zhao et al., 2024c) addresses quantization and model layer partitioning for heterogeneous clusters, incorporating a Cost Model and an Indicator Generator to optimize bit-width assignment and layer partitioning through integer linear programming, enhancing quantization for complex computational setups. Activation-aware Weight Quantization (AWQ) (Lin et al., 2024c) is a hardware-friendly, low-bit, weight-only quantization method for on-device LLMs that preserves key weights based on activation distribution, not weight distribution, to minimize quantization loss. AWQ uses per-channel scaling to optimize quantization error reduction without relying on backpropagation, thus maintaining LLMs’ generalization across different domains and modalities without overfitting.
Regarding the quantization evaluations of memory usage, MobileAIBench (Murthy et al., 2024) evaluates the impact of different quantization levels, including 3-bit, 4-bit, 8-bit, and 16-bit, on large language and multimodal models on iPhone 14. Firstly, it shows that quantization can significantly reduce disk usage while maintaining performance, e.g., Llama 2 7B’s disk usage decreases from 13 GB at 16-bit to 3.6 GB at 4-bit, while the F1 score in HotpotQA (Yang et al., 2018) only drops from 0.21 to 0.198. Secondly, while 4-bit quantization generally preserves effective performance, reducing to 3-bit significantly impairs model accuracy. For instance, Mobile-VLM-1.7B (Wu et al., 2024c) shows a marked decrease in accuracy on the VQA-v2 (Goyal et al., 2017) benchmark, dropping from 0.622 at 16-bit to 0.607 at 4-bit, and plummeting to 0.076 at 3-bit. This highlights that while more aggressive quantization levels can yield greater memory savings, they may drastically reduce model accuracy. Rahman et al. (2023) evaluates MobileBERT (Sun et al., 2020) on Raspberry Pi 3B edge devices using TensorFlow-Lite at 8-bit, 16-bit, and 32-bit quantizations. The evaluation uses crafted tweet data and a reputation polarity classification task. Results show 32-bit consumes 74.9 W/sample with 1.66s latency, while 8-bit reduces consumption to 47.49 W/sample and latency to 1.06s. Accuracy is nearly identical, at 0.685 for 32-bit and 0.684 for 8-bit. Memory usage is 64.1% for 32-bit and 66.1% for 8-bit. Quantization significantly reduces latency and energy consumption by approximately 1.5 times while maintaining performance levels. However, memory usage is not optimized due to the need for dequantization during inference. MobileLLM (Liu et al., 2024h) applies quantization to its MobileLLM and MobileLLM-LS models, which have 125M and 350M parameters respectively, resulting in a modest accuracy reduction of less than 0.5 points.
@@ 4.2.1. Memory Efficiency Optimization
Cache of MoE Experts. Beyond standard quantization, which reduces storage for all model parameters, another strategy involves caching a mixture of experts (MoE). Driven by the fact that memory storage is more cost-effective and scalable than computing capacity, the MoE architecture (Jacobs et al., 1991) boosts performance while minimizing computational costs by activating only portions of the LLM as needed. However, this approach incurs significant memory overhead, making it impractical for edge device memory constraints. For example, Switch Transformers (Fedus et al., 2022), with 32 experts per layer, require 54GBs of memory for inference, exceeding the capacity of most edge devices. Yi et al. (2023) notes that in the Switch Transformers model, although most of the model weights (86.5%) are attributed to experts, these weights account for only a small fraction (26.4%) of the computations. To address this, EdgeMoE (Yi et al., 2023) introduces a method where experts are loaded into an expert memory buffer only when activated, achieving approximately $3\times$ memory savings compared to the baseline where all weights are held in memory.
KV Cache Compression. When serving LMs for inference, using a KV cache is common practice to avoid intensive recalculations and speed up generation (Pope et al., 2023). However, cache memory consumption escalates with model size and sequence length, posing a challenge for edge devices. To manage this, offloading techniques transfer KV caches to CPU memory or storage (Sheng et al., 2023; Aminabadi et al., 2022), although this can introduce significant overhead due to the switching costs between GPUs and CPUs. An alternative approach, token dropping, compresses the cache size by retaining only the most important tokens, based on the commonly low attention scores (Zhang et al., 2024g; Liu et al., 2024b; Ge et al., 2024). Yet, such methods struggle with complex tasks requiring long responses or chain-of-thought (CoT) reasoning, as performance declines at high compression levels due to increased estimation errors—the disparity between the original and compressed KV values. To address these issues, GEAR (Kang et al., 2024) enhances KV cache quantization by integrating error-reduction techniques. This approach includes: (i) quantizing the majority of caches of similar magnitudes to ultra-low precision, (ii) using a low-rank matrix to efficiently approximate the quantization residuals, and (iii) employing a sparse matrix for a negligible ratio of entries of large magnitudes to correct individual errors from outliers. This composite approximation strategy separates the coherent from the incoherent parts of the approximation error: the low-rank matrix captures the majority of the coherent basis of the quantization error, while the sparse matrix addresses the inconsistencies present in individual outliers. This optimization enables near-lossless KV cache compression, achieving a peak memory reduction of up to $2.29\times$. Dynamic Memory Compression (DMC) (Nawrot et al., 2024) adaptively compresses the KV cache by choosing either to add the current key and value representations directly to the cache or to blend them with the top item in the cache using a weighted average. Evaluations on Llama 2 (Touvron et al., 2023b) models with 7B, 13B, and 70B parameters demonstrate that DMC enables LLMs to maintain downstream performance on benchmarks such as MMLU (Hendrycks et al., 2021) and HumanEval (Chen et al., 2021b) close to that of the original LLM. For instance, DMC Llama 2 7B with a $4\times$ compression ratio achieves an MMLU accuracy of 43.9, nearly matching the original model’s 44.6. Besides, Transformer-Lite (Li et al., 2024f) addresses a specific issue where model inputs and outputs redundantly store the KV cache twice, significantly increasing memory usage despite the content being largely identical. To optimize this, they refine KV cache storage by allocating a suitably large tensor for each input KV cache based on the maximum sequence length required for the inference task. Sub-tensors are then generated from this main tensor at varying address offsets, serving as model input and output KV caches. This approach allows the new KV caches to be written directly to the correct locations during model inference, thereby eliminating the need for additional copying processes. LLMaaS (Yin et al., 2024) introduces the concept of a Large Language Model as a Service on mobile devices, enabling LLMs to manage all device applications. LLMaaS proposes LLMS, which uses chunk-wise KV cache compression and swapping to facilitate context switching across apps within a limited memory budget. By dividing the KV cache into independently compressed and swapped chunks, LLMS optimally balances device memory use and I/O bandwidth, outperforming token-level or context-level management.

### v2

@@ 4.1.5. SLM Applications in Mobile-device
The use of cloud-based LLMs on devices raises privacy concerns and their large size limits real-time responses in urgent scenarios such as medical emergencies. To overcome these issues, researchers are creating smaller, domain-specific models (SLMs) that offer accurate results and suit mobile use. This subsection discusses SLM applications on mobile devices, focusing on three aspects: (1) software API calls, (2) mobile control, and (3) basic NLP applications.
SLM for Tool Learning. Integrating LLMs with APIs enhances capabilities but incurs high training costs, prompting a shift to smaller, task-specific models that cut costs but risk errors. In response, Octopus (Chen et al., 2024a) uses a diverse dataset from over 30K APIs and curriculum learning (Liu et al., 2024e) to improve API function accuracy. This method boosts API performance in models like Codellama-7b (Roziere et al., 2023) and Google’s Gemma series (Team et al., 2024a). PhoneLM-1.5B-Call (Yi et al., 2024) is fine-tuned on DroidCall (Xie et al., 2024c) datasets and achieves comparable performance compared to GPT-4o-mini (OpenAI, 2024a). $\alpha$-UMI (Shen et al., 2024a) employs SLMs as planners, callers, and summarizers within multi-agent systems, outperforming a single LLM in tool uses.
@@
SLM for Mobile Control. LM agents facilitate user-device interactions through taps, gestures, and text, automating tasks and enhancing user hands-free convenience. Unlike traditional developer-based approaches that require extensive developer effort to design interfaces and translate commands into API calls, LMs offer scalable automation via GUI-based text contents. MobileAgent (Ding, 2024) integrates instructions and Standard Operating Procedures (SOP) to improve SLMs for mobile control. As shown in Figure 16, it processes goals (e.g., booking a dental appointment) by analyzing screens, queries, prior actions, and UI elements, forming prompts to generate outputs and execute actions (e.g., selecting text, XPath). Fine-tuning Qwen-7B (Bai et al., 2023) on AIA medical data, it outperforms GPT-4 (Achiam et al., 2023) on AitW (Rawles et al., 2024), a key mobile benchmark, without extra inference costs. Carreira et al. (2023) run a small offline model on mobile devices, fine-tuned with ChatGPT-3.5 data, enabling tasks like calls and web searches. RedPajama-INCITE-Chat-3B-v1 Computer (2023), selected for its size and chat features, uses native code and quantization, performing well on 6GB and 4GB Android devices.
@@
AutoDroid (Wen et al., 2024) improves Android app interactions via GUI automation. Figure 17 shows LLM-powered task automation (e.g., setting a laundry reminder for Aug 17) in four steps: (1) click ’New Event’, (2) enter ’laundry’ in ’Title’, (3) click ’Save’, (4) finish. Using Vicuna-7B and app-specific knowledge, AutoDroid generates privacy-filtered prompts for tasks. On its DroidTask benchmark, it outperforms GPT-3.5 (34.7%) and GPT-4 (54.5%) with 57.7% accuracy. M4 (composable mobile foundation model) (Yuan et al., 2024c) introduces a 9.2B parameter model (7.5GB memory) for mobile AI tasks, managed by the OS and hardware. Currently limited to high-end devices, its applicability will expand with increasing mobile memory/storage. These works highlight customizing smaller, domain-specific SLMs to address memory limits while preserving functionality in mobile environments.
@@
Figure 18. Overview of fine-tuning SLMs with synthesized and user data (Qin et al., 2023). Data Synthesis generates semantically similar text via prompts, creating dialogue sets. Data Selection processes user data, tags domains, and calculates metrics (EOE, DSS, IDD). Selected data fine-tunes SLMs with user annotations. The iterative framework refines SLMs through continuous data generation and selection.
@@
SLM for Basic NLP Applications on Devices Performing basic NLP tasks such as text rewriting directly on the device can enable personalization while ensuring privacy. The sparse annotation on devices is a challenge. Qin et al. (2023) utilizes self-supervised data selection and synthesis for on-device fine-tuning, leveraging sparse annotations and limited storage effectively. This approach, demonstrated in Figure 18, employs the Llama-3B model (Touvron et al., 2023a) and the LoRA fine-tuning method (Hu et al., 2021), enhancing personalization by efficiently managing data through metrics including embedding entropy and domain-specific scores. In mobile text rewriting, Zhu et al. (2023b) train the compact Palm 2-XXS model (Anil et al., 2023) using data generated by the larger Palm 2-L to ensure user privacy and accommodate device constraints. On its new benchmark, MESSAGEREWRITEEVAL, Palm 2-XXS achieves a BLEU score of 34.59, outperforming LLaMa-7B (16.65) (Touvron et al., 2023a). Tests on the Samsung S23 show lower latency (29 tokens/s) than a 4-bit LLaMa-7B on a MacBook M1 Pro (18-22 tokens/s), proving its mobile efficiency for text rewriting.
@@
Insights: We draw several key insights from the development of task-specific SLMs: There is considerable potential in enhancing the efficiency and effectiveness of small models by integrating self-adaptive techniques with further fine-tuning and optimization of RAG-based methods. The growing relevance of SLMs in coding highlights their cost-effectiveness and efficiency as alternatives to LLMs, providing quick processing and easy fine-tuning for specialized tasks; while LLMs handle complex tasks well, SLMs, optimized and fine-tuned on specific data, are increasingly essential in resource-limited settings. SLMs significantly enhance recommendation systems due to their robust generalization, reasoning abilities, and in-context learning, addressing key challenges such as cold-start problems and distribution biases. They support long-term planning, replace traditional encoders, and use parallel low-rank parameters to inject personalized user knowledge effectively. SLMs play a crucial role in web search such as document encoding, text reordering, and query rewriting, often outperforming LLMs through techniques such as supervised fine-tuning, soft prompt tuning, unsupervised contrastive loss, and reinforcement learning, thereby enhancing adaptability and efficiency. SLMs are utilized on mobile devices primarily for privacy and memory constraints, with applications in API calls and mobile control; they are typically developed by generating data with LLMs and fine-tuning with SLMs, or by using local SLMs to handle privacy with LLMs boosting performance, and their training involves innovative techniques like learning from data streams and managing non-IID time series data.
4.2. SLM Deployment on Mobile and Edge Devices
4.2.1. Memory Efficiency Optimization
Compression on model parameters. Quantization, a common method for deploying SLMs, lowers numerical precision to significantly save memory while preserving accuracy. We detail quantization strategies in Sections 2.3.3 and 3.5, focusing here on representative works for edge devices. EDGE-LLM (Yu et al., 2024c) adapts LLMs for edge devices using a Layer-wise Unified Compression (LUC) method that combines layer-specific pruning and quantization to reduce computational demands and an Adaptive Layer Tuning and Voting scheme to optimize memory use while ensuring performance. Meanwhile, LLM-PQ (Zhao et al., 2024c) addresses quantization and model layer partitioning for heterogeneous clusters, incorporating a Cost Model and an Indicator Generator to optimize bit-width assignment and layer partitioning through integer linear programming, enhancing quantization for complex computational setups. Activation-aware Weight Quantization (AWQ) (Lin et al., 2024c) is a hardware-friendly, low-bit, weight-only quantization method for on-device LLMs, preserving essential weights based on activation distribution to minimize quantization loss. MoE-I2 (Yang et al., 2024f) prune less important experts and applies low-rank decomposition to the remaining experts to further optimize efficiency.
@@ 4.2.1. Memory Efficiency Optimization
AWQ (Lin et al., 2024c) Preserve key weights based on activation distribution.
MoE-I2 (Yang et al., 2024f) Prune less important experts in MoE.
@@
Transformer-Lite (Li et al., 2024g) Optimize KV cache to reduce redundancy and memory use.
@@
Runtime Efficiency Optimization mllm-NPU (Xu et al., 2024d) On-device NPU (neural processing units) to reduce prefill latency.
COST-EFF (Shen et al., 2022) Distill a multi-exit model from the original PLM.
@@
EdgeMoE (Yi et al., 2023) Predict expert needs, boosting inference speed and reducing latency.
@@
Cache of MoE Experts. Beyond standard quantization, which reduces storage for all model parameters, another strategy involves caching a mixture of experts (MoE). Driven by the fact that memory storage is more cost-effective and scalable than computing capacity, the MoE architecture (Jacobs et al., 1991) boosts performance while minimizing computational costs by activating only portions of the LLM as needed. However, this approach incurs significant memory overhead, making it impractical for edge device memory constraints. For example, Switch Transformers (Fedus et al., 2022), with 32 experts per layer, require 54GBs of memory for inference, exceeding the capacity of most edge devices. Yi et al. (2023) notes that in the Switch Transformers model, although most of the model weights (86.5%) are attributed to experts, these weights account for only a small fraction (26.4%) of the computations. To address this, EdgeMoE (Yi et al., 2023) introduces a method where experts are loaded into an expert memory buffer only when activated, achieving approximately 3×3$\times$3 × memory savings compared to the baseline where all weights are held in memory.
KV Cache Compression. When serving LMs for inference, using a KV cache is common practice to avoid intensive recalculations and speed up generation (Pope et al., 2023). However, cache memory consumption escalates with model size and sequence length, posing a challenge for edge devices. To manage this, offloading techniques transfer KV caches to CPU memory (Sheng et al., 2023; Aminabadi et al., 2022), although this can introduce significant overhead due to the switching costs between GPUs and CPUs. Token dropping compresses cache size by keeping only key tokens, often identified by low attention scores (Zhang et al., 2024g; Liu et al., 2024b; Ge et al., 2024). However, this method struggles with complex tasks, especially at high compression levels, due to increased estimation errors in compressed KV values. GEAR (Kang et al., 2024) addresses these issues by enhancing KV cache quantization with error-reduction techniques, including: (i) quantizing caches of similar magnitudes to ultra-low precision, (ii) using a low-rank matrix for efficient quantization residual approximation, and (iii) employing a sparse matrix for correcting outliers. This approach separates coherent from incoherent approximation errors, enabling near-lossless KV cache compression and achieving up to 2.29×2.29$\times$2.29 × peak memory reduction.
Besides, Dynamic Memory Compression (DMC) (Nawrot et al., 2024) adaptively compresses the KV cache by either adding new key and value representations directly or blending them with the top cache item using a weighted average. Transformer-Lite (Li et al., 2024g) tackles the redundancy of storing the KV cache twice in model inputs and outputs, which increases memory use. It optimizes storage by allocating a large tensor based on the maximum sequence length needed for inference. Sub-tensors are then created from this main tensor at different address offsets to serve as input and output KV caches, allowing direct writing to the correct locations during inference and removing extra copying steps. LLMaaS (Yin et al., 2024) introduces LLM as a Service for mobile devices, managing all apps through LLMS. This system uses chunk-wise KV cache compression and swapping, enabling efficient context switching within memory constraints. By segmenting the KV cache into independently compressed and swapped chunks, LLMS balances memory use and I/O bandwidth better than token-level or context-level management.